OpenStudy (anonymous):

Can someone help me simplify and check : 4/r+3+3/r+2=-4/r^2+5r+6

OpenStudy (mertsj):

\[\frac{4}{r+3}+\frac{3}{r+2}=\frac{-4}{(r+3)(r+2)}\]

OpenStudy (mertsj):

Multiply every term by (r+3)(r+2)